OpenStudy (calculusxy):

MEDAL! Find the ratios of volume of cylinder A to cylinder B. a) Cylinder A has the same radius but twice the height of cylinder B. b) Cylinder A has the same height but twice the radius of cylinder B.

OpenStudy (calculusxy):

@satellite73 @jim_thompson5910 @amistre64

OpenStudy (anonymous):

do you have an idea of what it is?

jimthompson5910 (jim_thompson5910):

Let's make up some numbers that fit the description in part (a) " Cylinder A has the same radius but twice the height of cylinder B." so let's say Cylinder A has radius r = 2, height h = 8 Cylinder B has radius r = 2, height h = 4 what is the volume of each cylinder?

OpenStudy (calculusxy):

So the formula to find the volume of a cylinder is V=Bh where B is the area of the base. Area of circle = \[\pi \times r^2\]

OpenStudy (calculusxy):

The radius of cylinder A is 2 so \[\pi \times 2^2\] is about 12.56 12.56 times height = 100.48

OpenStudy (calculusxy):

The radius of cylinder B is 2, but this time the height is different so: 12.56 times 4 = 50.24

OpenStudy (calculusxy):

This seems to be that the volume of cylinder A is twice as much as the volume of cylinder B.

jimthompson5910 (jim_thompson5910):

very good

OpenStudy (calculusxy):

So is it 2:1?

jimthompson5910 (jim_thompson5910):

Let's prove that using variables instead of numbers. So we can show that it is true in general no matter what r or h is Let r be the radius of both cylinders. Let h be the height of cylinder B. Cylinder A is twice as high, so its height is 2h

jimthompson5910 (jim_thompson5910):

Volume of Cylinder A V1 = pi*r^2*(2h) Volume of Cylinder B V2 = pi*r^2*(2h) divide the two volumes (V1/V2)